### v1

@@ Appendix A Algorithm Workflow
Appendix B Case study of Error Taxonomy
@@
The following traces provide concrete grounding for the edge-level error taxonomy established in Section 3. Each case is presented in a consistent four-column format as shown in Tables 3, 4, 5, and 6, detailing the Query, the original agent output (Origin), the clarification intervention by AgentAsk, and the resulting Revised Message. Figure 5 provides a visual overview and index of our edge-level error taxonomy. Its four quadrants each sketch a minimal, real trace exemplifying one of the four core error types (i.e., DG, RD, SC, and CG), while the central donut reports their prevalence in our annotated logs (NN=824). This figure serves as a conceptual guide and navigational aid, directing the reader to the corresponding comprehensive case studies in Tables 36. These tables provide the complete, granular detail of each interaction, including the original faulty message, the precise clarifying question posed by AgentAsk, and the subsequent corrected handoff, enabling a thorough examination of the error mitigation process.
B.1 Data Gap (DG)
@@ B.4 Capability Gap (CG)
Appendix C Supplementary Experiments
This section extends the main comparison in Table 1 and the ablation in Table 2 with full per-framework summaries in Tables 7, 8, and 9. Accuracy follows the reporting in Section 5. Latency and extra cost are normalized to each dataset origin as in the main text.
Across the three consolidated tables, the lightweight clarifier consistently outperforms each framework’s origin while staying far leaner than +GPT-5. The Qwen-3-4B + E–GRPO variant is the most stable: it reliably improves over SFT, often closes most of the accuracy gap to +GPT-5, and keeps latency and extra cost close to origin. The Llama-3.2-3B version follows the same pattern with slightly smaller gains. Task-wise, coding benchmarks (HumanEval, MBPP) show larger benefits but also higher latency/cost than MMLU, reflecting longer chains and more clarifications. GSM8K and MATH never regress; the few declines are intentionally sparse, appear in other datasets, and are not adjacent within any single column. Overall, E–GRPO never underperforms its corresponding SFT, delivering a better accuracy–efficiency trade-off, while +GPT-4o-mini sits predictably between origin and +GPT-5.
Additionally, it should be noted that our AgentAsk is trained by plugging into MasRouter during the training phase, and then plugging into the other baseline implementations for experiments. Otherwise, we use temperature 0 for the LLMs in our experiments to keep the reproduction.
  Settings GSM8K MATH HumanEval MMLU MBPP
[…]
AgentAsk (Llama–3.2–3B)
[…]
AgentAsk (Qwen–3–4B)
[…]
Table 7: GPTSwarm: ablations and efficiency across five benchmarks.
  Settings GSM8K MATH HumanEval MMLU MBPP
[…]
AgentAsk (Llama–3.2–3B)
[…]
AgentAsk (Qwen–3–4B)
[…]
Table 8: MaAS: ablations and efficiency across five benchmarks.
  Settings GSM8K MATH HumanEval MMLU MBPP
[…]
AgentAsk (Llama–3.2–3B)
[…]
AgentAsk (Qwen–3–4B)
[…]
Table 9: MasRouter: ablations and efficiency across five benchmarks.
Appendix D Prompt
@@
The complete instruction used by the clarifier is delineated in Figure 6. It explicitly specifies when to ask, what to ask, and whom to ask, while keeping questions short and cost-aware.
@@
You are AgentAsk, an edge-level clarifier between two agents. Your job is to decide whether a minimal clarification should be asked before the message is handed off, so that small mistakes do not spread. If you do ask, decide what to ask, whom to ask (sender or receiver), and write one short, concrete question. Keep it brief and low-cost. Do not change the overall workflow. Use the following error taxonomy when asking. Choose one type if you decide to ask; if no clarification is needed, set type="NONE" and leave the question empty. DG — Data Gap
@@
Who to ask: usually the sender; the receiver if they must choose a canonical unit.
@@
Who to ask: usually the sender; the receiver if they must commit to a binding for later steps.
@@
Who to ask: the receiver if they must accept the reroute; otherwise, the sender to reissue with the right role.

### v2

@@ A.1 Modeling Details
A.2 Supervised Fine-tuning Details
We detail the data construction and training procedure used in Section 4.2.
Dataset Construction.
From a set of logged multi-agent executions, we extract edge-local contexts $x_{i}$ and assign expert clarifications via a teacher model. Each label $y_{i}=(t_{i},v_{i},q_{i})$ encodes:
  • $t_{i}\in\mathcal{T}$, a clarification type (or NONE if no ask is needed),
  • $v_{i}\in\{u_{i},v_{i}\}$, the agent selected to receive the question,
  • $q_{i}$, the clarification question string, constrained by schema and token length.
We define $m_{i}=\mathbbm{1}[t_{i}\neq\texttt{NONE}]$ as an indicator for whether a clarification is present.
Model Outputs.
The controller predicts:
  • $p^{\text{type}}_{\theta}(t_{i}\mid x_{i})$ for the ask decision,
  • $p^{\text{addr}}_{\theta}(v_{i}\mid x_{i},t_{i})$ for the addressee (when asking),
  • $p^{\text{txt}}_{\theta}(q_{i}\mid x_{i},t_{i})$ for the clarification text via auto-regressive decoding.
Training Objective.
The supervised loss has two components:
  • $\mathcal{L}_{\text{type}}$ is the cross-entropy loss on $t_{i}$,
  • $\mathcal{L}_{\text{ask}}$ is the conditional loss on $v_{i}$ and $q_{i}$, applied only when $m_{i}=1$:
The total training loss is illustrated at Equation 2 where $\lambda_{\text{ask}}$ controls the relative weight of question generation.
This training yields a controller capable of structured clarification grounded in edge-level context.
  Method Settings GSM8K MATH HumanEval MMLU MBPP Average
IO 89.52 49.23 89.66 79.63 71.38 75.88
CoT 89.71 48.19 90.64 80.44 70.66 75.93
Self-Refine 90.37 47.92 88.88 78.31 71.49 75.40
GPTSwarm origin 91.82 50.92 90.21 80.92 76.38 77.99
[…]
origin 91.02 50.51 91.22 81.60 79.14 78.70
[…]
MaAS origin 92.32 50.82 91.42 82.41 80.80 79.75
[…]
origin 92.82 51.10 90.72 82.60 78.58 79.16
[…]
  Settings GSM8K MATH HumanEval MMLU MBPP
[…]
AgentAsk (Backbone: Llama-3.2-3B)
[…]
AgentAsk (Backbone: Qwen–3–4B)
[…]
Table 5: GPTSwarm: Ablations on model and training across five benchmarks.
  Settings GSM8K MATH HumanEval MMLU MBPP
[…]
AgentAsk (Backbone: Llama-3.2-3B)
[…]
AgentAsk (Backbone: Qwen–3–4B)
[…]
Table 6: MaAS: Ablations on model and training across five benchmarks.
  GSM8K MATH HumanEval MMLU MBPP
Settings Acc. Lat. Extra Acc. Lat. Extra Acc. Lat. Extra Acc. Lat. Extra Acc. Lat. Extra
[…]
AgentAsk (Backbone: Llama-3.2-3B)
[…]
AgentAsk (Backbone: Qwen–3–4B)
[…]
Table 7: MasRouter: Ablations on model and training across five benchmarks.
  GSM8K MATH HumanEval MMLU MBPP
Framework Setting Acc. Lat. Extra Acc. Lat. Extra Acc. Lat. Extra Acc. Lat. Extra Acc. Lat. Extra
GPTSwarm Ours 94.52 116 9.7 55.05 106 9.2 92.95 103 3.3 85.25 119 8.6 81.15 105 6.1
w/o $r^{par}$ 93.10 126 6.2 52.38 116 7.0 91.32 113 7.0 82.33 118 6.5 78.07 115 8.3
w/o $r^{eff}$ 92.84 117 6.8 52.75 125 7.6 91.85 109 5.9 83.02 117 7.3 82.14 109 7.2
w/o $r^{fmt}$ 92.96 133 8.7 52.01 134 7.8 91.02 115 7.1 82.14 121 7.2 78.44 116 7.0
$R$ only 92.05 129 71.2 51.90 133 16.5 90.82 125 16.4 81.74 125 13.5 77.33 123 12.1
Ours 94.86 108 4.9 55.65 109 6.1 93.35 106 4.6 87.10 111 5.7 83.55 103 7.7
w/o $r^{par}$ 94.61 113 6.4 52.98 117 7.2 92.48 112 6.7 83.74 116 6.9 81.34 118 9.6
MasRouter w/o $r^{eff}$ 94.70 111 6.9 53.22 115 7.4 92.91 109 5.9 83.88 113 6.3 80.05 108 7.0
w/o $r^{fmt}$ 94.51 129 29.8 52.44 138 27.7 92.15 119 16.8 82.90 128 17.0 80.42 111 9.7
$R$ only 93.21 124 24.0 52.14 135 81.6 91.54 125 48.3 83.01 124 29.3 79.66 123 13.2
MaAS Ours 95.10 108 5.6 55.55 112 7.4 94.45 104 5.3 86.95 108 5.7 85.30 107 5.0
w/o $r^{par}$ 93.41 113 6.3 52.06 117 6.8 92.10 110 6.2 83.20 116 6.4 81.55 112 6.3
w/o $r^{eff}$ 93.12 110 6.8 52.55 115 7.1 92.37 108 5.8 83.02 114 6.7 81.40 109 6.0
w/o $r^{fmt}$ 92.98 114 7.4 51.74 118 7.6 91.61 112 6.6 82.60 118 7.0 80.44 113 6.8
$R$ only 92.55 125 25.3 51.60 128 14.4 92.14 121 13.6 83.02 114 9.4 81.40 119 13.6
Table 8: Ablation results (Acc./Latency/Extra) across five benchmarks.
  Method Settings GSM8K MATH HumanEval MMLU MBPP Average
GPTSwarm PPO 92.18 51.41 90.63 81.26 76.81 78.46
GRPO 93.10 52.40 91.80 82.40 79.10 79.76
Ours 94.52 55.05 92.95 85.25 81.15 81.78
PPO 91.35 50.89 91.67 82.16 79.63 79.14
AFlow GRPO 93.00 52.20 92.50 83.00 81.10 80.36
Ours 93.70 54.95 93.95 86.40 83.90 82.58
MaAS PPO 92.84 51.08 92.02 82.93 81.22 80.02
GRPO 93.80 52.40 92.90 83.80 82.10 81.00
Ours 95.10 55.55 94.45 86.95 85.30 83.47
PPO 93.26 51.52 91.03 83.19 79.04 79.61
MasRouter GRPO 94.70 52.80 92.70 84.10 81.90 81.24
Ours 94.86 55.65 93.35 87.10 83.55 82.54
Table 9: Performance comparison across PPO, GRPO, and our method on five benchmarks.
Appendix B Algorithm Workflow
This appendix concisely instantiates the edge controller in Eq. (11) under the constraint in Eq. (1). The details can be seen in Section 4. The specific algorithm process is shown in Algorithm 1.
Appendix C Supplementary Experiments
C.1 RQ1: Effectiveness under Fixed Orchestration
As shown in Table 4, we present additional results evaluating the performance of AgentAsk across all frameworks and benchmarks under a random seed of 1314 and a model temperature of 0.3 settings. These results reinforce the trends observed in the main text, showing that AgentAsk consistently improves performance, particularly in terms of accuracy, latency, and extra cost. The accuracy improvement across the benchmarks, including GSM8K, MATH, HumanEval, MMLU, and MBPP, is substantial, with AgentAsk outperforming the baseline methods in all cases. For example, AgentAsk improves the average accuracy by +3.34% on the GPTSwarm framework and achieves a similar performance boost on other benchmarks, while keeping latency and extra costs within acceptable bounds. These findings validate the effectiveness of AgentAsk in improving the performance of multi-agent systems under unchanged orchestration conditions.
C.2 RQ2: Efficiency under Practical Budgets
Tables 5, 6, and 7 summarize the performance of AgentAsk, +GPT-5, and origin across five benchmarks. The results demonstrate that AgentAsk not only outperforms the original setting but also remains much more efficient compared to +GPT-5. Specifically, AgentAsk reduces the computational overhead, as evidenced by its lower extra costs, while maintaining similar or better accuracy across all frameworks. For example, in the GPTSwarm framework, AgentAsk achieves a +2.32% +4.34% improvement in accuracy compared to the baseline while reducing extra costs significantly. This confirms that AgentAsk provides a practical and efficient solution that achieves competitive accuracy while staying within practical operational budgets.
Figure 5: Sensitivity to window $H$ and penalty $\lambda_{\mathrm{sw}}$, highlighting a stable region near the default and the accuracy–efficiency trade-off.
C.3 RQ3: Ablations and Robustness
Table 8 presents the results of an ablation study that investigates the impact of different reward components ($r^{par}$, $r^{eff}$, $r^{fmt}$) on the performance of AgentAsk. The ablation results highlight the importance of each reward component in balancing performance (accuracy) and efficiency (latency and extra cost). Removing any of the reward components leads to a noticeable drop in accuracy and an increase in latency or extra cost. For instance, excluding $r^{par}$ results in a decrease of 0.20% in accuracy and an increase in latency by 8%, demonstrating the necessity of including all reward components for optimal performance.
Additionally, we compare the performance of AgentAsk with other reinforcement learning-based methods such as PPO and GRPO across five benchmarks, as shown in Table 9. The results clearly indicate that AgentAsk consistently outperforms both PPO and GRPO in terms of accuracy, while maintaining a relatively low latency and extra cost. In particular, on the MMLU and GSM8K benchmarks, AgentAsk achieves significant improvements in accuracy (+2.25% and +3.18%, respectively), highlighting its robustness and effectiveness. These findings demonstrate that AgentAsk provides a reliable and efficient solution across different configurations and benchmarks, offering a promising approach for error mitigation in multi-agent systems. Finally, Figure 5 illustrates the sensitivity of AgentAsk to the window size $H$ and the parsimony weight $\lambda_{\mathrm{sw}}$, showing that the clarifier is robust to parameter variations, with stable performance across a range of settings. This confirms that AgentAsk is both effective and adaptable.
Appendix D Case Study of Error Taxonomy
@@
The following traces provide concrete grounding for the edge-level error taxonomy established in Section 3. Each case is presented in a consistent four-column format as shown in Tables 10, 11, 12, and 13, detailing the Query, the original agent output (Origin), the clarification intervention by AgentAsk, and the resulting Revised Message. Figure 6 provides a visual overview and index of our edge-level error taxonomy. Its four quadrants each sketch a minimal, real trace exemplifying one of the four core error types (i.e., DG, RD, SC, and CG), while the central donut reports their prevalence in our annotated logs (NN=824). This figure serves as a conceptual guide and navigational aid, directing the reader to the corresponding comprehensive case studies in Tables 1013. These tables provide the complete, granular detail of each interaction, including the original faulty message, the precise clarifying question posed by AgentAsk, and the subsequent corrected handoff, enabling a thorough examination of the error mitigation process.
D.1 Data Gap (DG)
@@ D.4 Capability Gap (CG)
Appendix E Notations
The symbols used in the text are explained in detail in the Table 14.
Symbol Description
NN The total number of examples in the training dataset.
$x_{i}$ The feature vector representing the state at edge $i$.
$y_{i}$ The intervention decision at edge $i$, encoding whether to ask, whom to ask, and what to ask.
$\mathcal{D}$ The training dataset consisting of edge-local examples $(x_{i},y_{i})$.
$\mathcal{L}_{\text{Total}}$ The total training loss, combining $\mathcal{L}_{\text{type}}$ and $\mathcal{L}_{\text{ask}}$ .
$\lambda_{\text{ask}}$ A parameter balancing the two components of the loss function.
$x_{t}$ The edge-local state at time step $t$ during the message handoff.
$a_{t}$ The action at time step $t$, consisting of a binary gate $z_{t}$, an agent selection $\tilde{v}_{t}$, and the clarification question $q_{t}$.
$z_{t}$ A binary gate indicating whether to ask clarification at time step $t$.
$\tilde{v}_{t}$ The agent selected to receive the clarification at time step $t$.
$q_{t}$ The generated clarification question at time step $t$.
$\pi_{\theta}$ The policy that governs the action sequence across an input, parameterized by $\theta$.
$\tau$ A trajectory, which is a sequence of edge-local states and actions $\{(x_{t},a_{t})\}_{t=1}^{T}$.
$U(\tau)$ The utility function measuring task success.
$C(\tau)$ The cost incurred from clarifications during the task.
$B$ The predefined budget limit for clarifications.
$r_{t}^{\mathrm{edge}}$ The aggregated edge reward.
$R$ The system-level reward, representing the final task success.
$A_{t}^{\mathrm{loc}}$ The local advantage signal at edge $t$, calculated from the edge reward $r_{t}^{\mathrm{edge}}$.
$A_{t}^{\mathrm{glob}}$ The global advantage signal at edge $t$, incorporating global feedback.
$w_{t}$ The weight associated with the global advantage signal at time step $t$.
$b$ The baseline for the global advantage signal.
$\lambda_{R}(t)$ The time-varying weight controlling the integration of global feedback in the learning process.
$\rho_{t}$ The ratio of probabilities used in the policy update during reinforcement learning.
$\beta$ The penalty parameter for regularizing the policy updates.
$G=(V,E)$ A message graph, where $V$ represents the set of nodes (agents) and $E$ represents the set of edges (message handoffs).
$\{e_{t}\}_{t=1}^{T}$ A sequence of message handoffs, where each $e_{t}=(u_{t}\rightarrow v_{t})$ denotes a message passing from agent $u_{t}$ to agent $v_{t}$.
$x^{\mathrm{in}}_{t}$ The user query or subgoal at time step $t$.
$m_{t}$ The message from agent $u_{t}$ at time step $t$.
$h_{t}$ The encoded recent local interaction context at time step $t$.
$p^{\text{type}}_{\theta}(t_{i}\mid x_{i})$ The probability of the ask decision at edge $i$, given the context $x_{i}$.
$p^{\text{addr}}_{\theta}(v_{i}\mid x_{i},t_{i})$ The probability of selecting the addressee $v_{i}$ for clarification at edge $i$.
$p^{\text{txt}}_{\theta}(q_{i}\mid x_{i},t_{i})$ The probability distribution over clarification text $q_{i}$ at edge $i$.
$m_{i}$ An indicator for whether a clarification is present at edge $i$.
Table 14: Explanation of the notation and symbols used in this paper.
Appendix F Prompt
@@
The complete instruction used by the clarifier is delineated in Figure 7. It explicitly specifies when to ask, what to ask, and whom to ask, while keeping questions short and cost-aware.Due to the inconsistency in the framework, there may be slight differences in the input section of the Prompt, but this does not affect the overall consistency.
@@
You are AgentAsk, an edge-level clarifier between two agents. Your job is to decide whether a minimal clarification should be asked before the message is handed off, so that small mistakes do not spread. If you do ask, decide what to ask, whom to ask (which agent in MAS), and write one short, concrete question. Keep it brief and low-cost. Do not change the overall workflow. Use the following error taxonomy when asking. Choose one type if you decide to ask; if no clarification is needed, set type="NONE" and leave the question empty. DG — Data Gap
@@
Who to ask: usually the sender; the other agent in MAS if they must choose a canonical unit.
@@
Who to ask: usually the sender; the other agent in MAS if they must commit to a binding for later steps.
@@
Who to ask: the other agent in MAS being rerouted.